OpenStudy (anonymous):

Find the particular antiderivative of the following derivative that satisfies the given condition C'(x)=4X^2-3X C(0)=1,000

OpenStudy (anonymous):

@jim_thompson5910 you know you want to help me! :)

jimthompson5910 (jim_thompson5910):

First integrate both sides with respect to x \[\large C^{\prime}(x) = 4x^2 - 3x\] \[\large \int C^{\prime}(x)dx = \int (4x^2 - 3x)dx\] \[\large C(x) = \frac{4}{3}x^3-\frac{3}{2}x^2+D\] Note: I used D as the constant since C is used in the function name

jimthompson5910 (jim_thompson5910):

now we use the fact that C(0) = 1000 to find D \[\large C(x) = \frac{4}{3}x^3-\frac{3}{2}x^2+D\] \[\large C(0) = \frac{4}{3}(0)^3-\frac{3}{2}(0)^2+D\] \[\large 1000 = \frac{4}{3}(0)^3-\frac{3}{2}(0)^2+D\] \[\large 1000 = \frac{4}{3}(0)-\frac{3}{2}(0)+D\] \[\large 1000 = 0-0+D\] \[\large 1000 = D\] \[\large D = 1000\] So the function C(x) is \[\large C(x) = \frac{4}{3}x^3-\frac{3}{2}x^2+1000\]

OpenStudy (anonymous):

How do you know how to integrate it? Is it different for each indefinite integral problem

jimthompson5910 (jim_thompson5910):

when it says "find the antiderivative", that's the essentially the same as saying "integrate"

jimthompson5910 (jim_thompson5910):

the antiderivative is more descriptive because it literally means "opposite of derivative", so you're undoing the derivative....which is exactly what an indefinite integral is

OpenStudy (anonymous):

Can I pick your brain in a different example? And see how that differs?

jimthompson5910 (jim_thompson5910):

So for example say we have f(x) = x^2 we derive it to get f ' (x) = 2x ------------- now if we integrate (aka take the antiderivative), we would get integral of ( f ' (x) ) = x^2 + C so we kinda undid what the derivative did to x^2 to make it 2x...but there's that constant added to it (since the derivative of x^2 and x^2 + 1 is the same)

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Find the indefinite integral 1/2+5x^6 all times (30x^5) dx

jimthompson5910 (jim_thompson5910):

so it's \[\large \left(\frac{1}{2}+5x^6\right)(30x^5)\] right?

jimthompson5910 (jim_thompson5910):

or is it \[\large \left(\frac{1}{2+5x^6}\right)(30x^5)\]

OpenStudy (anonymous):